- BML.C.11.CO.Ly0 HV channel driving high current on 29th July [e-log].
- HV channel tripped after reaching ~250uA, no current visible in the gas gaps. HV channel ramped down to 9300V.
- BML.B.09.PI.Ly0 HV channel driving high current on 29th July [e-log].
- BML1A09.PI.DP2.Ly0.DZ1 gas gap driving high current. HV channel ramped down to 9300V.
- BMS.A.04.CO.Ly0 HV channel driving high current on 29th July [e-log].
- HV channel tripped after reaching ~180uA. Gas gap current are relatively not high. Icalib increased from 179uA to 185uA. HV channel ramped to nominal HV.
- BML.B.15.CO.Ly1 HV channel driving high current on 29th July [e-log].
- HV channel tripped due to BML2C15.CO.DP2.Ly1.DZ1 gas gap driving high current. HV Vconfig set to 9200V.
- BMS.A.16.CO.Ly0 HV channel driving high current on 30th July [e-log].
- HV channel tripped due to BMS3A16.CO.Ly0.DZ2 and SPC4A16.CO.DP2.Ly0.DZ1 gas gaps driving high current. HV Vconfig set to 9200V.
- BML.C.01.CO.Ly1 HV channel driving high current on 30th July [e-log].
- HV channel tripped due to BML3C01.CO.DP1.Ly1.DZ2 and BML3C01.CO.DP2.Ly1.DZ2 gas gaps driving high current. HV channel ramped down to 9150V.
- BML.C.11.CO.Ly0 HV channel driving high current on 30th July [e-log].
- HV channel tripped due to BML3C01.CO.DP1.Ly1.DZ2 and BML3C01.CO.DP2.Ly1.DZ2 gas gaps driving high current. HV channel ramped down to 9150V.
- RPC FATAL in side A/C and INFRASTRUCTURE due to 4 boards that went into unplugged status on 31st July [e-log, e-log, e-log].
- Firstly masked, recovered the following day sending the reset command to the affected crates on the CAEN crate console. Recovered 4 A3512AP HV modules, 2 A3801 ADC modules. All mainframes were reset.
- BOS.A2.02.CO.Ly1 HV channel driving high current on 1st August [e-log].
- HV channel tripped due to BOS4A02.CO.DP2.Ly0.DZ1 gas gap on the other layer driving high current. HV channel ramped down to 8800V.
- BOS.C2.16.CO.Ly1 HV channel driving high current on 2nd August [e-log].
- Overcurrent already at 3660V, HV channel masked and turned off.
- RPC resynchs in multiple sectors due to the noise bursts, leading to multiple stopless removals and towers killed [e-log].
- BOL.C2.07.CO.Ly0 HV channel driving high current on 3rd August [e-log].
- HV channel ramped down to 7000V due all connected gas gaps starting to drive high current from around 16:50.